OpenStudy (bobobox):

Which of the following is the correct graph of the solution to the inequality −13 > −5x + 2 > −28? (1 point) number line with open dots at 3 and 6 and shading in between. number line with a closed dot at 6 with shading to the right and another closed dot at 3 with shading to the left. number line with an open dot at 6 with shading to the right and another open dot at 3 with shading to the left. number line with closed dots at 3 and 6 with shading in between.

Directrix (directrix):

@bobobox Take this: −13 > −5x + 2 > −28 Add -2 -2 -2 -2 -------------------------------- What do you get?

OpenStudy (bobobox):

That looks very intimidating @Directrix

Directrix (directrix):

-13 - 2 =what

OpenStudy (bobobox):

11

OpenStudy (bobobox):

or -15?

Directrix (directrix):

-13 - 2 = - 15

OpenStudy (bobobox):

okay, bare with me hear I'm clueless

Directrix (directrix):

What are these: 2 -2 = -28 - 2 =

OpenStudy (bobobox):

0 and -30?

Directrix (directrix):

Right.

OpenStudy (bobobox):

okay... I promise I will medal

Directrix (directrix):

−13 > −5x + 2 > −28 -2 -2 -2 ------------------------ -15 > -5x > -30 is what we have so far.

OpenStudy (bobobox):

okay

Directrix (directrix):

Now, it is time to divide by -5

Directrix (directrix):

-15 divided by - 5 = -5x divided by - 5 = -30 divided by -5 =

OpenStudy (bobobox):

-3 -x? -6

Directrix (directrix):

A negative divided by a negative is positive. You have sign-us trouble.

OpenStudy (bobobox):

I have algebra troubles not just with signs

Directrix (directrix):

Don't be negative. Just change the signs on your answers and post. Let's keep moving.

OpenStudy (bobobox):

3 x 6

Directrix (directrix):

Good. Now, these numbers are part of an inequality expression. When you divide an inequality by a negative, then the direction of the inequality reverses. So, our greater thans (>) are reversing to less thans (<) Hold on.

Directrix (directrix):

-15 > -5x > -30 becomes 3 < x < 6 That is what we have to graph or describe. Look at your options. Do you see anything with a 2 and a 6 in it.

Directrix (directrix):

We are looking for numbers between 3 and 6 but not including 3 and 6.

OpenStudy (texaschic101):

if there is no equal signs, then your circles are open. greater then (>) shaded to the right and less then (<) shaded to the left

OpenStudy (bobobox):

number three im not quite sure I see some with a line passsing by two but I just dont want to screw up

Directrix (directrix):

@bobo Let me look and see what I think.

OpenStudy (texaschic101):

3 < x < 6....break it down 3 < x....x is greater then 3 x < 6 .... x is less then 6 so we have x being greater then 3 and less then 6.....does that help ?

Directrix (directrix):

@bobobox I assume that the number line started out as the color brown/gold and that the graphs are in blue.

OpenStudy (bobobox):

number line graph itself is gold and the options are blue... im pretty sure but hey anythings poosible

Directrix (directrix):

So, look for the blue segment with the "empty" circles at each end.

OpenStudy (bobobox):

okay

OpenStudy (texaschic101):

the shading is blue

OpenStudy (bobobox):

yup

Directrix (directrix):

@bobobox Which one do you see that looks like that?

OpenStudy (bobobox):

the first option

OpenStudy (texaschic101):

yep....you got it

Directrix (directrix):

@bobobox Yes, that is correct.

OpenStudy (bobobox):

Yay I'm sorry is there any way you guys could help me with to more I will fan the both of you and write...those...thing.. what ever they are called

OpenStudy (bobobox):

um i didn't invite anyone

OpenStudy (texaschic101):

I am sorry...I will butt out now. I didn't mean to cause a fuss

OpenStudy (bobobox):

I dont care who helps me I just need help

OpenStudy (bobobox):

Um I know what an interloper is

Directrix (directrix):

@bobobox Post your next question in a new thread. This one has gotten long.